OpenStudy (anonymous):

f(x)=ln(9+8x) find f'(x). show your work please.

OpenStudy (anonymous):

\[f'(x)=\frac{1}{9+8x}\times \frac{d}{dx}(9+8x)\] \['f(x)=\frac{8}{9+8x}\]

OpenStudy (anonymous):

o i see what i was doing wrong. thanks satellite!